how do u solve y=x-1
2x-y=0

Answer by fanism(39)

y=x-1 -- eq1
2x-y=0 -- eq2
sub eq1 into eq2
2x - (x-1) = 0
2x - x + 1 = 0
x + 1 = 0
x = -1
put back into eq1
y = -1 - 1
y = -2